OpenStudy (anonymous):

NEED HELP! A lunar lander, Apollo 10, is making its decent to Moon Base WPHS. The lander descends slowly under the retro-thrust of its descent engine. The engine is cut off when the lander is 5.0m above the surface and has a downward velocity of 0.8 m/s. With the engine off, the lander is in free fall what is the speed of the lander just before it touches the surface? The acceleration due to gravity on the moon is 1.6 m/s^2.

OpenStudy (anonymous):

Use: v^2=u^2 + 2as

OpenStudy (anonymous):

u=0.8 m/s s=5.0m a=1.6m/s^2

OpenStudy (anonymous):

Now, find v

OpenStudy (anonymous):

I forgot to find v

OpenStudy (anonymous):

v^2=(0.8)^2 + 2*1.6 *5

OpenStudy (anonymous):

Just simplify

OpenStudy (anonymous):

I GOT 16.64 it looks wrong

OpenStudy (anonymous):

That is right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that my final answer or I have to do more work

OpenStudy (anonymous):

Thats it

OpenStudy (anonymous):

ok thanks for all your help

OpenStudy (anonymous):

WELCOME